a garden 7 m by 12 m will be expanded by planting a border of flowers. the border will be ofthe same width around the entire garden and has an area of 92 square meter.

Answer by Boreal(15235)

The 92 m^2 will be (7+x)(12+x) to accommodate the border
84+19x+x^2=92
x^2+19x-8=0
x= (1/2) -19+/- sqrt (361+32), and the minus root won't make sense.
x=(1/2) (-19)+19.8242=0.4121 m
7.4*12.4=91.76, so this is a reasonable answer.